# Look-Ahead Reasoning on Learning Platforms

### v1

@@ Abstract
On many learning platforms, the optimization criteria guiding model training reflect the priorities of the designer rather than those of the individuals they affect. Consequently, users may act strategically to obtain more favorable outcomes, effectively contesting the platform’s predictions. While past work has studied strategic user behavior on learning platforms, the focus has largely been on strategic responses to a deployed model, without considering the behavior of other users. In contrast, look-ahead reasoning takes into account that user actions are coupled, and—at scale—impact future predictions. Within this framework, we first formalize level-$k$ thinking, a concept from behavioral economics, where users aim to outsmart their peers by looking one step ahead. We show that, while convergence to an equilibrium is accelerated, the equilibrium remains the same, providing no benefit of higher-level reasoning for individuals in the long run. Then, we focus on collective reasoning, where users take coordinated actions by optimizing through their joint impact on the model. By contrasting collective with selfish behavior, we characterize the benefits and limits of coordination; a new notion of alignment between the learner’s and the users’ utilities emerges as a key concept. We discuss connections to several related mathematical frameworks, including strategic classification, performative prediction, and algorithmic collective action.
@@ 1.1 Our contributions
We study look-ahead reasoning on learning platforms—user behavior that anticipates the actions of others in the population—by characterizing how it reshapes learning dynamics and equilibria. We begin with selfish agents, who act independently but strategically account for the other agents’ responses to the deployment of a predictive model. We then turn to coordinated behavior, where agents act jointly and strategically leverage their influence on the learner.
@@
Thus, if the average agent utility and the average loss of the learner are orthogonal, there is no benefit to coordination. Similarly, when $u=c\cdot\ell$ for either $c>0$ or $c<0$, the benefit of coordination is zero. However, when there is the right amount of overlap between the objectives that the collective can exploit, coordination can lead to more favorable outcomes than selfish reasoning. For example, think of instance $z$ as a feature-label pair $(x,y)$ and consider a learner and a collective who both care about predictions $f_{\theta}(x)$: the learner cares about making them accurate, and the collective cares about them having certain favorable values. Then, modifying the labels $y$ can be an effective strategy for the collective to steer the predictive model $\theta$, something that individuals acting selfishly cannot achieve. This overlap in the two sides’ objectives enables a large benefit of coordination. We elaborate on this example later on.
In additional results, we study heterogeneous populations made up of selfish agents and collectives of varying size. The results shed light on the benefits and limitations of coordination on learning platforms. For example, we explain why bigger collectives do not always lead to a higher average utility for the collective. We also show how broader participation in the collective stabilizes learning dynamics.
@@ 1.2 Background and related work
Strategic classification (hardt16strategic; bruckner2011stackelberg) introduces a model to study strategic behavior in learning systems based on assumptions of individual rationality. It describes a population of agents best-responding to a decision rule by altering their features to achieve positive predictions, given a fixed decision rule. Several variations of this basic model have been studied (e.g., dong2018strategic; chen2020learning; bechavod2021information; ghalme2021strategic; jag21alt); see podimata2025incentive for a recent survey of this literature. All these works focus on studying how agents strategize against a fixed decision rule. Our work introduces a new dimension of reasoning to strategic classification, taking into account how individual agents’ actions are coupled and how this influences the predictive model the agents strategize against.
@@
A more recent literature on algorithmic collective action (hardt2023algorithmic) studies coordinated agent efforts with the goal of steering learning systems; see (baumann2024algorithmic; ben2024role; gauthier2025statisticalcollusioncollectiveslearning; sigg2024decline) for recent developments in this area, as well as related discussions of data leverage (vincent21dataleverage). From the perspective of our work, collective action is a type of look-ahead reasoning: agents plan through model updates under the assumption that they coordinate with other agents. In this work we study the tradeoffs and implications of coordination on the population’s utility, assuming agents are evaluated against the data they report. Relatedly, hardt2022power discuss how platforms can reduce risk by actively steering a population. Collective action reverses this perspective and investigates how the population can improve its utility by steering the learner. This perspective is related to (zrnic2021leads), who also deviate from the classical model of strategic classification and instead model the population as the leader in the Stackelberg game against the learning platform. Our framework aims to bridge strategic classification and algorithmic collective action to illuminate the benefits and limits of coordination.
@@ 4 Collective reasoning
Since $h^{\sharp}$ is globally optimal, it holds that $\mathrm{B}\geq 0$. How large $\mathrm{B}$ is depends on the goals pursued by the learner and the population, as characterized by $\ell$ and $u$, respectively. Through coordinated data modifications, the population can steer the model towards a common target. But to do so, they may have to deviate from the individually optimal strategy. Thus, what governs the benefit of coordination is the tradeoff between the return of steering the model to a better equilibrium and the cost of being evaluated against suboptimal data.
@@
There are two different mechanisms at play, depending on the sign of $c$. In an adversarial setting where $c>0$, the game between the learner and the collective is a zero-sum game. In this case the benefit of coordination is zero, as the cost of steering is equal to its return. When $c<0$, the platform and the agents pursue the same goal and the game becomes a potential game between the two. In this case the benefit of coordination is zero as selfish actions are simultaneously optimal for the collective and there is no reason to steer the model away form the selfish equilibrium. In general, however, the benefit of coordination can be arbitrarily large. Consider the following example.
@@
We have $\Phi=0$ in the case where the gradients are orthogonal and the two functions $u$ and $\ell$ are unrelated; in this case agents can optimize their utility independent of the model and there is no benefit to being able to steer the model. Further, in line with Proposition 4, when $u=c\cdot\ell$ for some constant $c$, we also have $\Phi=0$ because selfish actions are simultaneously collectively optimal and shifting the model away from $\theta^{*}$ would not benefit the collective. In particular, $\theta^{*}$ is a local optimum of the loss under $\mathcal{D}_{h^{*}}$ and thus the first term in the inner product becomes zero. We provide a simple example satisfying the assumptions of Theorem 6 with a non-trivial closed-form expression for $\Phi$ in Appendix B.
To see cases where $\Phi^{2}$ can be large, we look at the individual terms in its definition. The first term in the inner product that defines $\Phi$ describes the suboptimality of $\theta^{*}$ for the agents under $\mathcal{D}_{h^{*}}$. It is non-zero if the collective’s utility on $\mathcal{D}_{h^{*}}$ can be improved by moving the model away from $\theta^{*}$. The second term in the inner product captures the response of the learner to the collective strategy. For this term to be non-zero the data modification $\mathcal{D}_{h^{\sharp}}$ must make $\theta^{*}$ suboptimal. Consequently, $\Phi$ is large if both terms in the inner product are non-zero, and there is overlap in the direction of improvement for the learner and the collective. Moreover, what really matters is not just the raw gradient alignment but one filtered through the local curvature of the loss landscape. The directions that the learner finds “flat” (small Hessian eigenvalues) allow for more influence on the model through small data modifications, and thus they offer more leverage for the collective.
Although Example 5 does not satisfy the assumptions in Theorem 6, the alignment term provides the right intuition for why $\mathrm{B}$ is large: since the learner aims to accurately predict labels, any systematic label modification induces a model change. The collective can fully control the direction of these changes, i.e., the gradient of the loss, and align it with their objectives. The label modification strategy leverages this to increase the agents’ utility, providing non-zero return as long as $\theta^{*}$ is suboptimal for the collective under $\mathcal{D}_{h^{*}}$.
@@ 5 Heterogeneous populations
A perfectly coordinated population, or one that implements the optimal collective strategy, is unlikely to emerge in practice. In the following we consider some plausible deviations from the idealized collective studied in the previous section and discuss how this impacts agent utilities and outcomes. Unless stated otherwise, we assume the collective implements any fixed strategy $h$ (independent of $\theta$), which could be a simpler alternative to a potentially hard-to-implement optimal strategy $h^{\sharp}$.
5.1 Fixed strategies in the presence of non-strategic agents
@@
The result provides a condition for determining whether a strategy is worth scaling up or not, evaluated with respect to the equilibrium it induces. It is a direct consequence of the envelope theorem and the condition is again linked to a notion of alignment described by $\Psi$. Suppose $\mathbf{H}$ is positive definite; then, if $u=\ell$, $\frac{\partial U_{\alpha}}{\partial\alpha}<0$, and if $u=$-$\ell$, $\frac{\partial U_{\alpha}}{\partial\alpha}>0$. Note that the result holds for any fixed strategy $h$. We provide additional intuition and empirical insights into how $U_{\alpha}$ and $\Psi$ change with $\alpha$ in Section 6. In particular, with a concrete example, we illustrate that larger collectives are not always better off and we demonstrate how the sign change of $\Psi$ accurately characterizes the optimal collective size $\alpha\in(0,1)$ for a fixed strategy $h$.
@@
Consider the mixture model (3) with $h=h_{\alpha}^{\sharp}$. Then, the utility $U^{*}_{\alpha}$ achieved by implementing the optimal size-aware strategy $h^{\sharp}_{\alpha}$ at any scale satisfies
@@
Note that the derivative in the first term takes into account the dependence of the strategy on $\alpha$. Thus, the result says that reoptimizing the strategy as a function of collective size does not change whether scaling up is worth it or not. All that matters is the alignment of the pursued goals. The argument involves considering how the equilibrium changes after reoptimizing the strategy and evaluating this change against the overall change in the population.
@@ 6.2 Trade-offs in collective reasoning
Figure 3: Collective utility decreases with collective size in the zero-sum case. The collective implements the optimal size-aware strategy for $\lambda=0$ in a mixed population with non-strategic agents. Small collectives can realize large gains, but the influence on the learner impedes gains at larger sizes $\alpha$.
The cost of steering.
@@ 7 Conclusion
A central goal of our work was to provide a unifying framework to contrast selfish reasoning with collective reasoning on learning platforms. We hope this opens up new pathways to transfer insights from strategic classification to algorithmic collective action, and vice versa. Beyond this, there are several natural extensions of our work. One is investigating look-ahead reasoning under imperfect information. As common in economic models, we assumed agents determine their actions under perfect information. In our case, this depends on the learner’s loss function and the population’s data distribution. However, in practice this information needs to be estimated from finite data which raises questions of statistical complexity. A related question would be to study how model misspecifications and estimation errors impact outcomes in look-ahead reasoning.
@@ Acknowledgements
Celestine Mendler-Dünner acknowledges the financial support of the Hector Foundation. The work was conducted during an internship of Haiqing Zhu at the MPI for Intelligent Systems, Tübingen.
@@ A.2 Proof of Theorem 3
The claim in the theorem follows by combining Lemma 11 with Theorem 3.5 in perdomo20performativeprediction. The key step is to prove Lemma 11 below.
@@
where the first inequality follows from Lemma˜10. ∎
@@ A.4 Proof of ˜4
Therefore, since $u=c\cdot\ell$, we can conclude that $\mathbb{E}_{z\sim\mathcal{D}_{h^{*}}}\left[\nabla_{\theta}u(z,\theta^{*})\right]=c\cdot\mathbb{E}_{z\sim\mathcal{D}_{h^{*}}}\left[\nabla_{\theta}\ell(z,\theta^{*})\right]=0$. Then, by Theorem˜12 below, we have $\mathrm{B}=0$.
A.5 Generalized version of Theorem˜6
We first state a general version of Theorem˜6 without the linearity assumption from Assumption 1.
Theorem 12 (Generalization of Theorem 6).
Suppose $U(h)$ is $\gamma$-strongly concave. Then, we have
$0\leq\mathrm{B}\leq\frac{1}{2\gamma}\cdot\left\|\mathbb{E}_{z\sim\mathcal{D}_{h^{*}}}\left[\nabla_{h}\nabla_{\theta}\ell(z,\theta^{*})\right](\mathbf{H}^{\star})^{-1}\mathbb{E}_{z\sim\mathcal{D}_{h^{*}}}\left[\nabla_{\theta}u(z,\theta^{*})\right]\right\|_{2}^{2}.$
where $\theta^{*}$ denotes the stable point corresponding to the selfish strategy $h^{*}$ and $\theta^{\sharp}$ denotes the stable point corresponding to the strategy $h^{\sharp}=\operatorname*{arg\,max}_{h}U(h)$.
Proof.
For notational simplicity, we set $f(h,\theta):=\mathbb{E}_{z\sim\mathcal{D}_{h}}\left[u(z;\theta)\right]$ and $g(h,\theta):=\mathbb{E}_{z\sim\mathcal{D}_{h}}\left[\ell(z;\theta)\right]$. Recall that
$U(h)=f(h,\mathcal{A}(\mathcal{D}_{h}))=\mathbb{E}_{z\sim\mathcal{D}_{h}}\left[u(z,\mathcal{A}(\mathcal{D}_{h}))\right],$
where the first argument in $f$ only applies in distribution that the distribution is taken against and the second argument is corresponding to the second argument of $u$. Then, by the implicit function theorem we have
$\nabla_{h}U(h)=\nabla_{h}f(h,\mathcal{A}(\mathcal{D}_{h}))=\nabla_{1}f(h,\mathcal{A}(\mathcal{D}_{h}))-\nabla^{2}_{1,2}g(h,\mathcal{A}(\mathcal{D}_{h}))\left[\nabla^{2}_{2,2}g(h,\mathcal{A}(\mathcal{D}_{h}))\right]^{-1}\nabla_{2}f(h,\mathcal{A}(\mathcal{D}_{h})),$
where $\nabla_{1}$ and $\nabla_{2}$ denote the gradient operator on the first/second argument of the function. For the equilibrium $(h^{*},\theta^{*})$, we must have $\nabla_{1}f(h,\mathcal{A}(\mathcal{D}_{h}))=0$. Hence, by the PL-inequality, we obtain that
$f\left(h^{\sharp}\right)-f\left(h^{*}\right)\leq\frac{1}{2\gamma}\cdot\left\|\nabla^{2}_{1,2}g(h^{*},\mathcal{A}(\mathcal{D}_{h^{*}}))\left[\nabla^{2}_{2,2}g(h^{*},\mathcal{A}(\mathcal{D}_{h^{*}}))\right]^{-1}\nabla_{2}f(h^{*},\mathcal{A}(\mathcal{D}_{h^{*}}))\right\|_{2}^{2}.$
A.6 Proof of Theorem˜6

### v2

@@ Abstract
On many learning platforms, the optimization criteria guiding model training reflect the priorities of the designer rather than those of the individuals they affect. Consequently, users may act strategically to obtain more favorable outcomes. While past work has studied strategic user behavior on learning platforms, the focus has largely been on strategic responses to a deployed model, without considering the behavior of other users. In contrast, look-ahead reasoning takes into account that user actions are coupled, and—at scale—impact future predictions. Within this framework, we first formalize level-$k$ thinking, a concept from behavioral economics, where users aim to outsmart their peers by looking one step ahead. We show that, while convergence to an equilibrium is accelerated, the equilibrium remains the same, providing no benefit of higher-level reasoning for individuals in the long run. Then, we focus on collective reasoning, where users take coordinated actions by optimizing through their joint impact on the model. By contrasting collective with selfish behavior, we characterize the benefits and limits of coordination; a new notion of alignment between the learner’s and the users’ utilities emerges as a key concept. Look-ahead reasoning can be seen as a generalization of algorithmic collective action; we thus offer the first results characterizing the utility trade-offs of coordination when contesting algorithmic systems.
@@ 1.1 Our contributions
We characterize how look-ahead reasoning—user behavior that anticipates the actions of others in the population—reshapes learning dynamics and equilibria on learning platforms. We begin with selfish agents, who act independently but strategically account for the other agents’ responses to the deployment of a predictive model. We then turn to coordinated behavior, where agents act jointly and strategically leverage their influence on the learner. Finally, contrasting the two settings allows us to characterize the benefits and limitations of coordination on learning platforms.
@@
Thus, if the average agent utility and the average loss of the learner are orthogonal, there is no benefit to coordination. Similarly, when $u=c\cdot\ell$ for either $c>0$ or $c<0$, the benefit of coordination is zero. However, when there is the right amount of overlap between the objectives that the collective can exploit, coordination can lead to more favorable outcomes than selfish reasoning. For example, think of instance $z$ as a feature-label pair $(x,y)$ and consider a learner and a collective who both care about predictions $f_{\theta}(x)$: the learner cares about making them accurate, and the collective cares about them having certain favorable values. Then, modifying the labels $y$ can be an effective strategy for the collective to steer the predictive model $\theta$ through the learner’s response, something that individuals reasoning selfishly cannot achieve. This overlap in the two sides’ objectives enables a large benefit of coordination. We elaborate on this example later on.
In additional results, we study heterogeneous populations made up of selfish agents and collectives of varying size. The results explain why larger collectives do not always lead to a higher utility for participating agents. They also show how broader participation in the collective stabilizes learning dynamics.
@@ 1.2 Background and related work
Strategic classification (hardt16strategic; bruckner2011stackelberg) introduces a model to study strategic behavior in learning systems based on assumptions of individual rationality. It describes a population of agents best-responding to a decision rule by altering their features to achieve positive predictions, given a fixed decision rule. Several variations of this basic model have been studied (e.g., dong2018strategic; chen2020learning; bechavod2021information; ghalme2021strategic; jag21alt; levanon22ageneralSC); see podimata2025incentive for a recent survey of this literature. All these works focus on studying how agents strategize against a fixed decision rule. Our work introduces a new dimension of reasoning to strategic classification, taking into account how individual agents’ actions are coupled and how this influences the predictive model the agents strategize against.
@@
A more recent literature on algorithmic collective action (hardt2023algorithmic) studies coordinated agent efforts with the goal of steering learning systems; see (baumann2024algorithmic; ben2024role; gauthier2025statisticalcollusioncollectiveslearning; sigg2024decline) for recent developments in this area, as well as related discussions of data leverage (vincent21dataleverage) and protective optimization technologies (kulynych20POT). From the perspective of our work, collective action is a type of look-ahead reasoning: agents plan through model updates under the assumption that they coordinate with other agents. In this work we study the tradeoffs and implications of coordination on the utility of agents participating in the collective. Relatedly, hardt2022power discuss how platforms can reduce risk by actively steering a population. Collective action reverses this perspective and investigates how the population can improve its utility by steering the learner. This perspective is related to (zrnic2021leads), who also deviate from the classical model of strategic classification and instead model the population as the leader in the Stackelberg game against the learning platform. Our framework aims to bridge strategic classification and algorithmic collective action to illuminate the benefits and limits of coordination.
@@ 4 Collective reasoning
Since $h^{\sharp}$ is the globally optimal strategy for the population, it holds that $\mathrm{B}\geq 0$. How large $\mathrm{B}$ is depends on the goals pursued by the learner and the population, as characterized by $\ell$ and $u$, respectively. Through coordinated data modifications, the population can steer the model towards a common target. But to do so, they may have to deviate from the individually optimal strategy. Thus, what governs the benefit of coordination is the tradeoff between the return of steering the model to a better equilibrium and the cost of being evaluated against the modified data.
@@
There are two different mechanisms at play, depending on the sign of $c$. In an adversarial setting where $c>0$, the game between the learner and the collective is a zero-sum game. In this case the benefit of coordination is zero, as the cost of steering is equal to its return. When $c<0$, the platform and the agents pursue the same goal and the game becomes a potential game between the two. In this case the benefit of coordination is zero as selfish actions are simultaneously optimal for the collective and there is no benefit to steering the model away from the selfish equilibrium. In general, however, the benefit of coordination can be arbitrarily large. Consider the following example.
@@
We have $\Phi=0$ in the case where the gradients are orthogonal and the two functions $u$ and $\ell$ are unrelated; in this case agents can optimize their utility independent of the model and there is no benefit to being able to steer the model. Further, in line with Proposition 4, when $u=c\cdot\ell$ for some constant $c$, we have $\Phi=0$ because selfish actions are simultaneously collectively optimal and shifting the model away from $\theta^{*}$ would not benefit the collective. In particular, $\theta^{*}$ is a local optimum of the loss under $\mathcal{D}_{h^{*}}$ and thus the first term in the inner product becomes zero.
To see cases where $\Phi^{2}$ can be large, we look at the individual terms in its definition. The first term in the inner product that defines $\Phi$ describes the suboptimality of $\theta^{*}$ for the agents under $\mathcal{D}_{h^{*}}$. It is non-zero if the collective’s utility on $\mathcal{D}_{h^{*}}$ can be improved by moving the model away from $\theta^{*}$. The second term in the inner product captures the response of the learner to the collective strategy. For this term to be non-zero the data modification $\mathcal{D}_{h^{\sharp}}$ must make $\theta^{*}$ suboptimal. Consequently, $\Phi$ is large if both terms in the inner product are non-zero, and there is overlap in the direction of improvement for the learner and the collective. Moreover, what really matters is not just the raw gradient alignment but one filtered through the local curvature of the loss landscape. The directions that the learner finds “flat” (small Hessian eigenvalues) allow for more influence on the model through small data modifications, and thus they offer more leverage for the collective. We provide a simple example satisfying the assumptions of Theorem 5 with a non-trivial closed-form expression for $\Phi$ in Appendix B.
Let us revisit Example 1. Although the example does not satisfy the assumptions in Theorem 5, the alignment term provides the right intuition for why $\mathrm{B}$ is large: since the learner aims to accurately predict labels, any systematic label modification induces a model change. The collective can fully control the direction of these changes, i.e., the gradient of the loss, and align it with their objectives. The label modification strategy leverages this to increase the agents’ utility, providing non-zero return as long as $\theta^{*}$ is suboptimal for the collective under $\mathcal{D}_{h^{*}}$.
@@ 5 Heterogeneous populations
A perfectly coordinated population where every agent participates, or one that implements the optimal collective strategy, is unlikely to emerge in practice. In the following we consider some plausible deviations from the idealized collective studied in the previous section and discuss how this impacts agent utilities and outcomes. Unless stated otherwise, we assume the collective implements any fixed strategy $h$ (independent of $\theta$), which could be a simpler alternative to a potentially hard-to-implement optimal strategy $h^{\sharp}$.
5.1 Scaling strategies in the presence of non-strategic agents
@@
The result provides a condition for determining whether a strategy is worth scaling up or not, evaluated with respect to the equilibrium it induces. It is a direct consequence of the envelope theorem and the condition is again linked to a notion of alignment described by $\Psi$. To provide intuition for the result we consider some special cases. Suppose $\mathbf{H}$ is positive semidefinite, which happens when $\ell$ is convex; then, if $u=\ell$, $\Psi\geq 0$, and if $u=$-$\ell$, $\Psi\leq 0$. This means that, under convex losses, when considering the utility of participating agents, scaling up a fixed strategy is always harmful in a zero-sum game, and it is always beneficial when the learner and the population optimize the same target. Note that the result holds for any fixed strategy $h$. We provide additional intuition and empirical insights into how $U_{\alpha}$ and $\Psi$ change with $\alpha$ in Section 6. In particular, with a concrete example, we illustrate why larger $\alpha$ can be a disadvantage for the collective, and how the sign change of $\Psi$ accurately characterizes the optimal collective size $\alpha\in(0,1)$ for a fixed strategy $h$.
@@
Consider the mixture model (3) with a collective of size $\alpha$ implementing the optimal size-aware strategy $h=h_{\alpha}^{\sharp}$. Then, the average collective utility $U^{*}_{\alpha}$ achieved by implementing $h^{\sharp}_{\alpha}$ satisfies
@@
Note that the derivative in the first term takes into account the dependence of the strategy on $\alpha$. Thus, the result implies that reoptimizing the strategy as a function of collective size does not change whether scaling up is worth it or not. All that matters is the alignment of the pursued goals. The argument involves considering how the equilibrium changes after reoptimizing the strategy for the new size $\alpha$ and evaluating this change against the overall change in the population. A consequence is that, in the case of a zero-sum game and convex losses, the maximum utility a collective can achieve decreases with size.
@@ 6.1 Retraining dynamics under level-$k$ thinking
Figure 3: Collective utility decreases with collective size in the zero-sum case. The collective implements the optimal size-aware strategy for $\lambda=0$ in a mixed population with non-strategic agents. Small collectives can realize large gains, but the response by the learner impedes gains at larger sizes $\alpha$.
@@ 6.2 Trade-offs in collective reasoning
Cost of steering and why larger collectives can be worse off.
@@ 7 Conclusion
A central goal of our work was to provide a unifying framework to contrast selfish reasoning with collective reasoning when contesting machine learning predictions. There are several natural extensions of our work. One is investigating look-ahead reasoning under imperfect information. As common in economic models, we assumed agents determine their actions under perfect information. In our case, this concerns knowledge about the learner’s loss function and the population’s data distribution. However, in practice this information needs to be estimated from finite data which raises questions of statistical complexity. A related question would be to study how model misspecifications, estimation errors, and imperfect coordination impact outcomes in look-ahead reasoning.
More broadly, we hope our work can support recent discussions concerning data poisoning and multi-party adversarial attacks (e.g., nestaas2025adversarial), and open up new pathways to transfer insights from strategic classification to algorithmic collective action, and vice versa.
@@ Acknowledgements
Celestine Mendler-Dünner acknowledges the financial support of the Hector Foundation. The work was conducted during an internship of Haiqing Zhu at the Max Planck Institute for Intelligent Systems, Tübingen.
@@ A.1 Auxiliary results
Below we state a bound on the benefit of coordination that does not assume linearity (Assumption 1).
Theorem 5.
Let each strategy $h(\eta)$ be represented by a parameter vector $\eta\in\mathbb{R}^{d}$. Suppose $U(h(\eta))$ is $\gamma$-strongly concave on $\eta$. Then, we have
$0\leq\mathrm{B}\leq\frac{1}{2\gamma}\cdot\left\|\nabla_{\eta}\nabla_{\theta}\mathbb{E}_{z\sim\mathcal{D}_{h^{*}}}\left[\ell(z,\theta^{*})\right](\mathbf{H}^{\star})^{-1}\mathbb{E}_{z\sim\mathcal{D}_{h^{*}}}\left[\nabla_{\theta}u(z,\theta^{*})\right]\right\|_{2}^{2}.$
with $\mathbf{H}^{\star}:=\mathbb{E}_{z\in\mathcal{D}_{h^{*}}}\left[\nabla_{\theta,\theta}^{2}\ell(z,\theta^{*})\right]$ and $\theta^{*}$ denoting the stable point corresponding to the selfish strategy $h^{*}$.
@@ A.2 Proof of Theorem 3
The key step is to prove Lemma 10 below. The claim in the theorem follows by combining Lemma 10 with Theorem 3.5 in perdomo20performativeprediction.
@@
where the first inequality follows from Lemma˜9. ∎
@@ A.4 Proof of Proposition 4
Therefore, since $u=c\cdot\ell$, we can conclude that $\mathbb{E}_{z\sim\mathcal{D}_{h^{*}}}\left[\nabla_{\theta}u(z,\theta^{*})\right]=c\cdot\mathbb{E}_{z\sim\mathcal{D}_{h^{*}}}\left[\nabla_{\theta}\ell(z,\theta^{*})\right]=0$. Then, by ˜5, we have $\mathrm{B}=0$.
A.5 Proof of Theorem˜5